OpenStudy (anonymous):

write an exponential equation from two points. (3, 1.5) (4, 15)

OpenStudy (anonymous):

So you have \[y = Ae^{Rx}\] Taking each x,y pair and plugging into that equation will give you two equations you can use to find A and R

OpenStudy (anonymous):

um the equation my teacher gave me was y=ab^x. I got the answer of y=.0015(10)^x and I was just checking to make sure it was correct. :0]

OpenStudy (anonymous):

Ok. Your b = My e^R. \[1.5 = ab^3\] \[15 = ab^4\] Dividing them we have \[\frac{15}{1.5} = \frac{ab^4}{ab^3} \rightarrow 10 = b\] Then plugging back in for b we have \[1.5 = a(10)^3\rightarrow 0.0015 = a\] So yeah \[y = .0015(10)^x\]

OpenStudy (anonymous):

WHOOP! Thanks a bunch Lego Man!! Your assistance is highly appreciated!!

OpenStudy (anonymous):

Anytime =)